Your sister just deposited $9,000 into an investment account. She believes that she will earn an annual return of 9.5 percent for the next 8 years. You believe that you will only be able to earn an annual return of 8.6 percent over the same period. How much more must you deposit today in order to have the same amount as your sister in 8 years?
Answer #1

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

=>

future value of sister deposit

= 9000 * (1+9.5%)^8

= 18601.82

future value of brother deposit

= 9000 * (1+8.6%)^8

= 17413.30

Difference in future value = 18601.82 - 17413.30

= 1188.52

Present value of difference

= 1188.52/(1+8.6%)^8

= 614.28

Brother must deposit 614.28 more to have same amount inn 8 years